OpenStudy (anonymous):

A gas is found to be 85.7 carbon and 14.3 hydrogen by mass. its molecular mass was found to be 98.1 g/mol using the method of this experiment. What is the molecular formula of this gas?

OpenStudy (anonymous):

Let there is 100 g of this gas. Mass of C: 85.7 g Mass of H: 14.3 g No. of mole of C: 85.7/12=7.14 No. of mole of H: 14.3/1=14.3 Mole ratio between C and H= 1:2 therefore, the empirical formula of this gas: CH2 molecular mass: 98.1=(12+1+1)xN N=7 So the molecular formula of this gas is C7H14.